# Quasinormal Modes and Stability Analysis of the JMN-1 Naked Singularity

### v1

@@ Abstract
In this paper, we perform a comprehensive analysis of the quasinormal modes in the Joshi-Malafarina-Narayan (JMN-1) naked singularity by investigating its response to linear perturbations, including scalar, electromagnetic, and gravitational perturbations. To analyze the stability of the JMN-1 naked singularity under axial perturbations, we compute the quasinormal mode frequencies using the Wentzel-Kramers-Brillouin method. The quasinormal mode frequencies provides information about the stability of spacetime, with the real part of the frequency determining the oscillation rate and the imaginary part governing the decay or growth of perturbations. Our results indicate that by imposing appropriate boundary conditions, we find that the JMN-1 spacetime remains dynamically stable under axial perturbations.
@@ I Introduction
Extending the OSD framework, Roger Penrose proposed the Cosmic Censorship Conjecture (CCC), which postulates that the final singularity resulting from gravitational collapse must always remain concealed within an event horizon, ensuring that the collapse invariably leads to black hole formation penrose . Despite extensive efforts to establish a rigorous mathematical formulation or proof for CCC, no conclusive resolution has been reached. This remains one of the most fundamental open problems in gravitational physics, with significant implications for black hole theory and astrophysical applications. Consequently, various studies have explored more physically realistic collapse scenarios within Einstein gravity, incorporating inhomogeneous matter distributions and nonzero pressure profiles joshi ; Joshi:2011zm ; Joshi:2011rlc ; Joshi:2024gog ; goswami ; Deshingkar:1998ge ; Jhingan:2014gpa ; Joshi:2013dva . These studies raise the important question of how to observationally distinguish naked singularities from conventional black hole spacetimes. Therefore, various observational characteristics of naked singularities have been investigated, including their shadows and accretion disk properties Joshi:2013dva ; Saurabh:2023otl ; gyulchev ; Tahelyani:2022uxw ; Kovacs:2010xm ; Guo:2020tgv ; Joshi:2020tlq ; Saurabh:2022jjv ; Solanki:2021mkt ; Bambhaniya:2021ybs ; Bambhaniya:2024lsc ; Vagnozzi:2022moj , periastron precession of relativistic orbits Bambhaniya:2019pbr ; Joshi:2019rdo ; Dey:2019fpv ; Bam2020 ; Bambhaniya:2025xmu ; Bambhaniya:2022xbz ; Bambhaniya:2021jum , pulsar timings Kalsariya:2024qyp , tidal force effects Madan:2022spd , energy extraction mechanisms Patel:2023efv ; Patel:2022jbk ; Acharya:2023vlv , etc.
@@
The study of QNMs in black hole spacetimes has been extensively explored Konoplya:2019 ; Bhattacharya:2023 ; Konoplya:2003 ; Matyjasek_2017 ; Konoplya:2023 ; Zhao:2024 ; Jusufi:2020 ; Abdalla:2006 , revealing fundamental aspects of black hole dynamics, the nature of event horizons, and even potential observational signatures in gravitational wave astronomy Kokkotas:1999bd ; Berti:2009kk . However, the investigation of QNMs in naked singularity spacetimes remains an active and relatively less explored area of research. As we have pointed out above, based on metric tests and shadow images, the JMN-1 and JNW naked singularities are the most promising candidates to mimic conventional black hole models EventHorizonTelescope:2022xqj ; Saurabh:2023otl ; gyulchev ; Vagnozzi:2022moj . The QNMs and stability of the JNW naked singularity have been explored in Chowdhury:2020 ; Stashko:2023ffs . On the other hand, the JMN-1 spacetime, as a well-defined naked singularity solution, presents an excellent opportunity for studying the behavior of perturbations in such a background geometry. Whether it exhibits stable oscillatory behavior or instability remains an important aspect to investigate.
The stability analysis via QNMs can reveal whether the JMN-1 solution is physically viable or if perturbations lead to growing modes frequencies that may ruled out such spacetimes. In this work, we analyze the QNM frequencies using the Wentzel-Kramers-Brillouin (WKB) method in the JMN-1 naked singularity spacetime by considering linear perturbation expansion of the metric and investigate the stability of this solution under scalar, electromagnetic, and gravitational linear perturbations.
We have the following arrangement of the paper. In section (II), we perform the analysis of perturbations in the JMN-1 spacetime. In section (III), we discuss the boundary conditions for the JMN-1 naked singularity spacetime. In section (IV), we carry out the analysis of WKB method to study the perturbations. In section (V), we study the frequencies of fundamental QNM. In section (VI), we discuss and conclude the results of this work. Throughout the paper, we use the metric sign as ,+,+,+-,+,+,+- , + , + , + and the units $G=c=1$.
@@ II Perturbations in JMN-1 Spacetime
The parameter $M_{0}$ is a dimensionless quantity that characterizes the compactness of the object and the intensity of anisotropic pressure effects. It must satisfy the constraint $50<M_{0}<4/$ 5 to ensure that the speed of sound remains below the speed of light. The effective sound speed of the equilibrium state, given by $c_{s}^{2}=p/\rho$, must remain subluminal, thus necessitating $M_{0}<4/5$ Joshi:2011zm . The Schwarzschild mass of the system is given by $M=\frac{1}{2}M_{0}R_{b}$, establishing a direct relationship between $M_{0}$ and the total gravitational mass. Physically, a higher $M_{0}$ results in stronger gravitational effects near the core, yielding a more compact structure. Notably, $R_{b}$ must not be less than $2.5M$ to satisfy the sound speed condition.
@@
The perturbed metric gives rise to the perturbed Christoffel symbols:
@@
Due to the spherical symmetry of the background spacetime, we can decompose the perturbations into odd and even types. In this work, we focus only on the axial (odd-parity) perturbations of the background metric. Note that here we assume that the anisotropic matter fluid constituting the JMN-1 naked singularity within $R_{b}$ does not contribute to axial (odd-parity) perturbations. In many cases, anisotropic matter fluids with only tangential pressure do not contribute to odd-parity perturbations Chowdhury:2020 ; C:2024cnk ; Zhang:2024bov ; Blazquez-Salcedo:2013ava ; Kobayashi:2012kh . The reason is that odd-parity perturbations involve shear like perturbations of the metric but do not involve scalar quantities like energy density or pressure directly. In such cases, the axial perturbations still follow a Regge-Wheeler (RW) like equation. However, some specific models of an anisotropic matter fluid do contribute to the axial perturbations, especially if there is shear viscosity or if the fluid stress-energy tensor has nonzero off-diagonal terms in an orthonormal frame. Since we consider JMN-1 as a spherical symmetric spacetime, the evolution of the axial perturbation for the asymptotically flat spacetime is governed by the following field equation:
@@
$\delta R_{\theta\phi}$ $=-\frac{\partial}{\partial t}\left[\left(-\frac{h_{0}}{2f(r)}\right)\left(\frac{\partial}{\partial\theta}(\sin\theta\,\partial_{\theta})P_{l}(\cos\theta)\right)\right]$
$\quad-\frac{\partial}{\partial r}\left[\frac{f(r)}{2}\left(\frac{\partial h_{1}}{\partial\theta}\right)(\sin\theta\,\partial_{\theta})P_{l}(\cos\theta)\right]=0.$ (15)
Define,
@@
$\frac{d^{2}\tilde{\psi}(r_{*})}{dr_{*}^{2}}+\left(\omega^{2}-V(r_{*})\right)\tilde{\psi}(r_{*})=0.$ (20)
@@ III Boundary Conditions
$\psi(r)\sim\frac{e^{i\omega r}}{r},$ (21)
with $\omega$ being the complex QNM frequency. At sufficiently large $r$, the perturbation equation stabilizes, ensuring that the outgoing wave condition, holds accurately.
$\frac{d\psi}{dr}\approx i\omega\psi,$ (22)
Studies of QNMs in Schwarzschild and Kerr spacetimes Kokkotas:1999bd ; Berti:2009kk have demonstrated that setting the outer boundary at moderate distances is sufficient to extract accurate mode frequencies, as further increases in $r$ do not significantly affect the results.
Unlike black hole spacetimes, where the inner boundary for QNM analysis is set at the event horizon with a purely ingoing wave condition, the JMN-1 naked singularity presents a unique challenge due to the absence of an event horizon. Since the central singularity at $r=0$ is globally naked, we must introduce an alternative inner boundary condition to ensure a well-posed perturbation problem. Therefore, we impose an inner boundary at $r=M_{0}R_{b}$. We take $M_{0}=0.7$ and $R_{b}=2.8571$, where matter distribution extends up to $R_{b}$, and beyond this radius, the spacetime is vacuum. In this matter distributed region ($r<R_{b}$), ingoing waves are expected to be absorbed by the singularity. Besides, the inner boundary can be considered up to some orders of the Planck length.
@@ IV WKB method
The parameter $s$ gives the spin-weight of the perturbing field. Since we are dealing with scalar, electromagnetic, and gravitational perturbations, we use $s=0$, $s=1$ and $s=2$, respectively. Unlike black holes, the JMN-1 naked singularity lacks an event horizon, making boundary conditions crucial: an inner boundary near $r=0$ and an outer boundary at large $r$, where outgoing waves define stability. The behavior of $Q(x)$ dictates whether perturbations decay or persist, making its precise formulation essential in analysis of the stability of the singularity under the given perturbation.
@@
Figure 2: Convergence of quasinormal mode frequencies with increasing WKB order for obertone number $n=0$, $l=2$ amd $s=2$. The red color corresponds for imaginary and the blue color corresponds to real QMN frequencies.
The presence of increasingly large terms in the expansion can lead to numerical instabilities, as observed in the $12^{th}$-order results. This highlights the trade-off between precision and practicality when using the WKB method. In our analysis, we find that the sixth and seventh orders provide the most reliable results before numerical errors start dominating. The convergence as also seen in Figure 2 as these orders aligns well with previous studies on WKB-based QNM computations, reinforcing their applicability in gravitational wave research. Moreover, the deviation at extremely high orders suggests that alternative resummation techniques, such as Padé approximants, could be explored to enhance the stability of WKB calculations further. These refinements may help in extending the method’s validity to more extreme astrophysical scenarios, including highly compact objects and beyond-GR modifications. In this paper, we calculate the QNMs using the publicly available Mathematica package Konoplya202X .
@@ V The QNMs frequencies
In this section, we present the results of our numerical calculations. Tables 1 to 5 display the real and imaginary parts of the QNM frequencies for three primary types of axial perturbations: scalar, electromagnetic, and gravitational. Scalar perturbations $(s=0)$ arise from massless scalar fields propagating in the given background spacetime. These perturbations do not carry spin and are often used to study the behavior of test fields in strong gravitational fields. Electromagnetic perturbations $(s=1)$ correspond to the perturbations of the Maxwell field in the given background. The QNMs associated with electromagnetic perturbations are particularly relevant in analyzing the stability of naked singularities in the presence of electromagnetic radiation. Finally, gravitational perturbations $(s=2)$ are associated with fluctuations in the metric tensor itself, representing the fundamental response of the spacetime geometry to small disturbances. The data from the tables show imaginary negative parts of the QNM frequencies across all cases, therefore, confirming the stability of the spacetime.
@@
0.1 $0.373619-0.088891i$ $0.582642-0.28129i$
0.3 $0.373657-0.0888999i$ $0.5827-0.281319i$
0.5 $0.373619-0.088891i$ $0.582642-0.28129i$
0.6 $0.373657-0.0888999i$ $0.5827-0.281319i$
0.7 $0.371286-0.0883358i$ $0.579003-0.279534i$
Table 3: Real and imaginary parts of the fundamental QNM frequencies in Scalar field perturbations ($s=0$) for $6^{th}$ order WKB approximation.
$M_{0}$ $l=2,n=0$ $l=3,n=1$
$\omega(WKB)$ $\omega(WKB)$
0.1 $0.483642-0.0967661i$ $0.660671-0.292288i$
0.3 $0.48369-0.0967758i$ $0.660737-0.292317i$
0.5 $0.483642-0.0967661i$ $0.660671-0.292288i$
0.6 $0.48369-0.0967758i$ $0.660737-0.292317i$
0.7 $0.480621-0.0961617i$ $0.656545-0.290462i$
@@
[…]
For gravitational (s=2), scalar (s=0), and electromagnetic (s=1) perturbations, the real parts of the QNMs remain nearly constant, with a slight decrease at $M_{0}=0.7$ as shown in tables 1,2, 3, and 4. The imaginary parts also show minimal variation, this indicates a stable decay rate across all cases. These trends suggest that the JMN1 spacetime maintains its stability without significant deviations in the QNM spectrum.
@@
When a system undergoes an initial perturbation, its response carries crucial information about its stability and underlying structure. In the analysis of the JMN-1 naked singularity, axial gravitational perturbations evolve over time, exhibiting a characteristic decay pattern. Initially, the perturbations undergo a phase of QNM ringing, where oscillations gradually decrease in amplitude. This stage would be followed by a decay phase, where the perturbations diminish at a slower rate. The rate at which perturbations dissipate depends strongly on the angular momentum mode $l$.

### v2

@@ Abstract
In this paper, we perform a comprehensive analysis of the quasinormal modes in an external geometry of the Joshi-Malafarina-Narayan (JMN-1) naked singularity by investigating its response to linear perturbations, including gravitational and electromagnetic perturbations. To analyze the stability of the JMN-1 naked singularity under axial perturbations, we compute the quasinormal mode frequencies using the Wentzel-Kramers-Brillouin method. The quasinormal mode frequencies provides information about the stability of spacetime, with the real part of the frequency determining the oscillation rate and the imaginary part governing the decay or growth of perturbations. Our results indicate that by imposing appropriate boundary conditions, we find that the background spacetime of JMN-1 naked singularity remains dynamically stable under axial perturbations.
@@ I Introduction
Extending the OSD framework, Roger Penrose proposed the Cosmic Censorship Conjecture (CCC), which postulates that the final singularity resulting from gravitational collapse must always remain concealed within an event horizon, ensuring that the collapse invariably leads to black hole formation penrose . Despite extensive efforts to establish a rigorous mathematical formulation or proof for CCC, no conclusive resolution has been reached. This remains one of the most fundamental open problems in gravitational physics, with significant implications for black hole theory and astrophysical applications. Consequently, various studies have explored more physically realistic collapse scenarios within Einstein gravity, incorporating inhomogeneous matter distributions and nonzero pressure profiles joshi ; Joshi:2011zm ; Joshi:2011rlc ; Joshi:2024gog ; goswami ; Deshingkar:1998ge ; Jhingan:2014gpa ; Joshi:2013dva . These studies raise the important question of how to observationally distinguish naked singularities from conventional black hole spacetimes. Therefore, various observational characteristics of naked singularities have been investigated, including their shadows and accretion disk properties Joshi:2013dva ; Saurabh:2023otl ; gyulchev ; Tahelyani:2022uxw ; Kovacs:2010xm ; Guo:2020tgv ; Joshi:2020tlq ; Saurabh:2022jjv ; Solanki:2021mkt ; Bambhaniya:2021ybs ; Bambhaniya:2024lsc ; Vagnozzi:2022moj ; Bambi:2019tjh , periastron precession of relativistic orbits Bambhaniya:2019pbr ; Joshi:2019rdo ; Dey:2019fpv ; Bam2020 ; Bambhaniya:2025xmu ; Bambhaniya:2022xbz ; Bambhaniya:2021jum , pulsar timings Kalsariya:2024qyp , tidal force effects Madan:2022spd , energy extraction mechanisms Patel:2023efv ; Patel:2022jbk ; Acharya:2023vlv , etc.
@@
The study of QNMs in black hole spacetimes has been extensively explored Konoplya:2019 ; Bhattacharya:2023 ; Konoplya:2003 ; Matyjasek_2017 ; Konoplya:2023 ; Zhao:2024 ; Jusufi:2020 ; Abdalla:2006 ; Ghosh:2023etd , revealing fundamental aspects of black hole dynamics, the nature of event horizons, and even potential observational signatures in gravitational wave astronomy Kokkotas:1999bd ; Berti:2009kk . However, the investigation of QNMs in naked singularity spacetimes remains an active and relatively less explored area of research. As we have pointed out above, based on metric tests and shadow images, the JMN-1 and JNW naked singularities are the most promising candidates to mimic conventional black hole models EventHorizonTelescope:2022xqj ; Saurabh:2023otl ; gyulchev ; Vagnozzi:2022moj . The QNMs and stability of the JNW naked singularity have been explored in Chowdhury:2020 ; Stashko:2023ffs . On the other hand, the JMN-1 spacetime, as a well-defined naked singularity solution, presents an excellent opportunity for studying the behavior of perturbations in such a background geometry. Whether it exhibits stable oscillatory behavior or instability remains an important aspect to investigate.
The stability analysis via QNMs can reveal whether the JMN-1 solution is physically viable or if perturbations lead to growing modes frequencies that may ruled out such spacetimes. In this work, we analyze QNM frequencies using the Wentzel-Kramers-Brillouin (WKB) method in the background metric of JMN-1 naked singularity by considering linear perturbation expansion of the metric and investigate the stability of this solution under gravitational and electromagnetic perturbations. In the stability analysis based on test fields, “scalar” refers to an external Klein–Gordon field propagating on a fixed background, treated analogously to electromagnetic or gravitational test fields. Such scalar field perturbations, by definition, do not appear in the axial (odd-parity) sector; instead, they are associated with the polar (even–parity) sector.
We have the following arrangement of the paper. In section (II), we perform the analysis of perturbations in an external spacetime of JMN-1 naked singularity. In section (III), we discuss the boundary conditions for the JMN-1 naked singularity spacetime. In section (IV), we carry out the analysis of WKB method to study the perturbations. In section (V), we study the frequencies of fundamental QNM. In section (VI), we discuss and conclude the results of this work. Throughout the paper, we use the metric sign as ,+,+,+-,+,+,+ and the units $G=c=1$.
@@ II Perturbations in JMN-1 Spacetime
The parameter $M_{0}$ is a dimensionless quantity that characterizes the compactness of the object and the intensity of anisotropic pressure effects. It must satisfy the constraint $0<M_{0}<4/5$ to ensure that the speed of sound remains below the speed of light. The effective sound speed of the equilibrium state, given by $c_{s}^{2}=p/\rho$, must remain subluminal, thus necessitating $M_{0}<4/5$ Joshi:2011zm . The Schwarzschild mass of the system is given by $M=\frac{1}{2}M_{0}R_{b}$, establishing a direct relationship between $M_{0}$ and the total gravitational mass. Physically, a higher $M_{0}$ results in stronger gravitational effects near the core, yielding a more compact structure. Notably, $R_{b}$ must not be less than $2.5M$ to satisfy the sound speed condition.
@@
It is important to emphasize that throughout this paper, our analysis is restricted to the exterior region of the JMN-1 spacetime, which ensures asymptotic flatness Bambhaniya:2019pbr . In particular, we consider this scenario to be analogous to the Schwarzschild black hole, where stability analyses are typically conducted in the external region extending from $2M$ to spatial infinity. In the black hole case, the ingoing plane waves are entirely absorbed by the event horizon located at the inner boundary. In contrast, for the JMN-1 naked singularity, perturbations may be partially scattered, reflected, or fully absorbed at the inner boundary or at the singularity. The behavior of these ingoing waves might depend on the specific matter profile. In this work, we focus solely on the external geometry of the JMN-1 naked singularity to avoid additional complexities. The study of perturbations within the boundary radius $R_{b}$ in the JMN-1 spacetime will be addressed separately in future work, as it requires alternative methods to investigate the stability of the interior region.
@@
In addition to gravitational perturbations, we consider electromagnetic field perturbations on the fixed background spacetime. For this field, we work with the vector potential $A_{\mu}$, which is decomposed as
$A_{\mu}=\bar{A}_{\mu}+\delta A_{\mu},$ (7)
where $\bar{A}_{\mu}$ is the background potential (vanishing in the case of an uncharged black hole spacetime), and $\delta A_{\mu}$ encodes the perturbations.
Now for the gravitational case, the perturbed metric gives rise to the perturbed Christoffel symbols:
@@
Due to the spherical symmetry of the background spacetime, we can decompose the perturbations into odd and even types.
In this work, we focus exclusively on the axial (odd-parity) perturbations of the background metric. We assume that the anisotropic matter fluid constituting the JMN-1 naked singularity within $R_{b}$ does not contribute to these perturbations. In many scenarios, anisotropic fluids with purely tangential pressure do not affect odd-parity perturbations Chowdhury:2020 ; C:2024cnk ; Zhang:2024bov ; Blazquez-Salcedo:2013ava ; Kobayashi:2012kh . This is because odd-parity perturbations involve shear-like distortions of the metric rather than scalar quantities such as energy density or pressure. In such cases, the axial perturbations obey a Regge-Wheeler (RW)-like equation. However, certain specific models of anisotropic fluids can contribute to axial perturbations, particularly when shear viscosity is present or when the stress-energy tensor exhibits nonzero off-diagonal components in an orthonormal frame. Since we consider JMN-1 as a spherical symmetric spacetime, the evolution of the axial perturbation for the asymptotically flat spacetime is governed by the following field equation:
@@
$\delta R_{\theta\phi}$ $=-\frac{\partial}{\partial t}\left[\left(-\frac{h_{0}}{2f(r)}\right)\frac{\partial}{\partial\theta}\left(\sin\theta\,\frac{\partial}{\partial\theta}P_{l}(\cos\theta)\right)\right]$
$\quad-\frac{\partial}{\partial r}\left[\frac{f(r)}{2}h_{1}\frac{\partial}{\partial\theta}\left(\sin\theta\,\frac{\partial}{\partial\theta}P_{l}(\cos\theta)\right)\right]=0$ (16)
Defining
@@
Similarly, the general relativistic wave equations for the electromagnetic ($A_{\mu}$) field can be written as follows:
$\frac{1}{\sqrt{-g}}\,\partial_{\mu}\left(F_{\rho\sigma}\,g^{\rho\nu}g^{\sigma\mu}\sqrt{-g}\right)$ $=0,$ (21)
where $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ is the electromagnetic field strength tensor. The effective potential for the electromagnetic ($s=1$) fields then takes the form
$V_{\text{eff}}(r)=f(r)\left[\frac{\ell(\ell+1)}{r^{2}}\right],$ (22)
where $\ell=s,s+1,s+2,\dots$ are the multipole numbers.
@@
$\frac{d^{2}\tilde{\psi}(r_{*})}{dr_{*}^{2}}+\left(\omega^{2}-V_{eff}(r_{*})\right)\tilde{\psi}(r_{*})=0.$ (23)
@@ III Boundary Conditions
$\Psi(r)\sim e^{-i\omega(t-r_{*})}$ (24)
At sufficiently large $r$, the perturbation equation stabilizes, ensuring that the outgoing wave condition is accurate. Studies of QNMs in Schwarzschild and Kerr spacetimes Kokkotas:1999bd ; Berti:2009kk have demonstrated that setting the outer boundary at moderate distances is sufficient to extract accurate mode frequencies, as further increases in $r$ do not significantly affect the results.
Unlike black hole spacetimes, where the inner boundary for QNM analysis is set at the event horizon with a purely ingoing wave condition, the JMN-1 naked singularity presents a unique challenge due to the absence of an event horizon. Since the central singularity at $r=0$ is globally naked, we must introduce an alternative inner boundary condition to ensure a well-posed perturbation problem. Therefore, we impose an inner boundary at $r=R_{b}$. Classically, it is expected that at the inner boundary, ingoing waves are fully absorbed by the central singularity. As a result, the wave function near this region takes the form
$\Psi(r)\sim e^{-i\omega(t+r_{*})},$ (25)
representing purely ingoing behavior. We take $M_{0}=0.7$ and $R_{b}=2.8571$, where matter distribution extends up to $R_{b}$, and beyond this radius, the spacetime is vacuum. The internal dynamics ($r<R_{b}$), where the spacetime is supported by an anisotropic fluid, play a crucial role in the complete perturbation analysis of the system. However, in this work our focus has been deliberately restricted to the exterior region ($r>R_{b}$) for the following reasons: (i) The interior JMN-1 spacetime ($r<R_{b}$) is already obtained as an equilibrium or stable configuration from gravitational collapse Joshi:2011zm , (ii) Our primary aim is to investigate the stability of the spacetime under axial perturbations in the vacuum domain, where the perturbation equations take a tractable Regge–Wheeler type form. Similarly, in the context of black holes, most of the literature (e.g., Schwarzschild, Kerr and Reissner–Nordström spacetime perturbations) has focused primarily on the exterior geometry, while the stability of the interior has largely been investigated using quantum gravity methods Boehmer:2008fz , (iii) The interior region involves non-vacuum anisotropic matter sources, which significantly complicates the perturbation analysis and requires a different methodology, typically involving coupled fluid–metric perturbations Andersson:2002jd ; Gundlach:1999bt . This separation of focus is a standard practice in perturbation theory, for example, in the case of the Sun and other stars, helioseismology studies often analyze oscillations by first treating the vacuum exterior separately ChristensenDalsgaard:2002 ; Kokkotas:1999bd . Motivated by this well established tradition, we have here confined ourselves to the exterior geometry of compact objects (JMN-1 naked singularity). Furhtermore, within the matter-distributed region ($r<R_{b}$), ingoing waves are expected to be partially absorbed by the singularity for a certain range of the parameter $M_{0}$ Acharya:2023vlv , or they may undergo partial reflection, producing echoes Chowdhury:2020 . We will address this issue in greater detail in a separate work. Moreover, the inner boundary may be considered down to scales of the order of the Planck length when analyzing the stability of the internal structure of this spacetime.
@@ IV WKB method
The parameter $s$ gives the spin weight of the perturbing field. Since we are dealing with electromagnetic, and gravitational perturbations, we use $s=1$ and $s=2$, respectively. The behavior of $Q(x)$ dictates whether perturbations decay or persist, making its precise formulation essential in analysis of the stability of the singularity under the given perturbation.
@@
Figure 2: Convergence of quasinormal mode frequencies with increasing WKB order for obertone number $n=0$, $l=2$ amd $s=2$. The red color corresponds to imaginary and the blue color corresponds to real QMN frequencies.
The presence of increasingly large terms in the expansion can lead to numerical instabilities, as observed in the $12^{th}$-order results. This highlights the trade-off between precision and practicality when using the WKB method. In our analysis, we find that the sixth and seventh orders provide the most reliable results before numerical errors start dominating. The convergence as also seen in Figure 2 as these orders aligns well with previous studies on WKB-based QNM computations Konoplya202X , reinforcing their applicability in gravitational wave research. Moreover, the deviation at extremely high orders suggests that alternative resummation techniques, such as Padé approximants, could be explored to enhance the stability of WKB calculations further. These refinements may help in extending the method’s validity to more extreme astrophysical scenarios, including highly compact objects and beyond-GR modifications. In this paper, we calculate the QNMs using the numerical WKB package Konoplya202X .
@@ V The QNMs frequencies
In this section, we present the results of our numerical calculations. Tables 1 to 4 display the real and imaginary parts of the QNM frequencies for the two primary types of axial perturbations: electromagnetic and gravitational. Electromagnetic perturbations $(s=1)$ correspond to the perturbations of the Maxwell field in the given background. Finally, gravitational perturbations $(s=2)$ are associated with fluctuations in the metric tensor itself, representing the fundamental response of the spacetime geometry to small disturbances. The data from the tables show imaginary negative parts of the QNM frequencies across all cases, therefore, confirming the stability of the spacetime.
@@
0.1 0.3736190.088891i0.373619-0.088891i 0.5826420.28129i0.582642-0.28129i
0.3 0.3736570.0888999i0.373657-0.0888999i 0.58270.281319i0.5827-0.281319i
0.5 0.3736190.088891i0.373619-0.088891i 0.5826420.28129i0.582642-0.28129i
0.6 0.3736570.0888999i0.373657-0.0888999i 0.58270.281319i0.5827-0.281319i
0.7 0.3712860.0883358i0.371286-0.0883358i 0.5790030.279534i0.579003-0.279534i
@@
9 0.37166 - 0.088915ii 0.0958424
8 0.371489 - 0.088917ii 0.00386661
7 0.37126 - 0.0882506ii 0.000123047
6 0.371286 - 0.0883358ii 0.0000732704
5 0.371172 - 0.088363ii 0.000120327
4 0.37122 - 0.0885673ii 0.000225863
3 0.370831 - 0.0886602ii 0.00248593
For gravitational (s=2), and electromagnetic (s=1) perturbations, the real parts of the QNMs remain nearly constant, with a slight decrease at $M_{0}=0.7$ as shown in tables 1,2, 3, and 4. The imaginary parts also show minimal variation, this indicates a stable decay rate across all cases. These trends suggest that the JMN1 spacetime maintains its stability without significant deviations in the QNM spectrum.
@@
When a system undergoes an initial perturbation, its response carries crucial information about its stability and underlying structure. In the analysis of the JMN-1 naked singularity, axial gravitational perturbations evolve over time, exhibiting a characteristic decay pattern. Although the time-dependent evolution of the perturbation is not explicitly depicted here, it typically reveals some important dynamical features. Initially, the perturbations undergo a phase of quasinormal ringing, where oscillations gradually decrease in amplitude. This stage is followed by a decay phase, where the perturbations diminish at a slower rate. The rate at which perturbations dissipate depends strongly on the angular momentum mode $l$.